# Fixed points of Lie group actions on moduli spaces A tale of two actions

### v1

@@ 1 Introduction
The aforementioned examples were found using the relationships between the various topological solitons and related objects. For instance, the examples of instantons and hyperbolic monopoles were found using their corresponding ADHM data—quaternionic matrices satisfying non-linear constraints. The topological solitons themselves do not fit into the framework studied in this paper. Indeed, the space of objects and the space of gauge transformations are infinite-dimensional. However, the related objects do fit in this framework and previous work of mine has used these related objects to identify novel examples of instantons, hyperbolic monopoles, and Euclidean monopoles [CDL+22, Lan24a, Lan24b].
@@
This work differs from the above in that we examine general Lie group actions instead of finite, abelian, or absolute Galois group actions. This choice of action allows us to study different kinds of symmetries. Moreover, by choosing Lie group actions, we can make use of smoothness to differentiate the equations of symmetry, providing a linear constraint, which greatly simplifies any potential non-linear constraints. Indeed, my previous work using these results have identified many novel topological solitons [CDL+22, Lan24a, Lan24b]. Additionally, this work differs as we do not restrict ourselves to a specific moduli space and group action.
In Section 2, we prove Theorem 1.1, which deals with compact Lie groups and first appeared in my thesis [Lan24b, Theorem 1.1.1 & Appendix A]. As compact Lie groups can be realized as matrix Lie groups, the Lie algebra homomorphisms give us representations of Lie algebras, so we reduce the problem to the realm of representation theory. We also prove Proposition 1.2, which allows us to differentiate the equations of symmetry when the symmetry group is one-dimensional but not necessarily compact. In Section 2.1, we use Proposition 1.2 to find symmetric points in an example. For in depth examples of the use of this result, see my thesis as well as my previous papers, where I identify novel examples of symmetric instantons as well as hyperbolic and Euclidean monopoles [CDL+22, Lan24a, Lan24b]. In Section 2.2, we show that the group of symmetries of an object is a closed Lie subgroup, which helps narrow down the kinds of symmetries that we need to study when looking at objects with differing continuous symmetries.
@@ 2 Fixed points
Theorem 1.1 requires compact groups. However, hope is not necessarily lost when you do not have compactness. Indeed, in previous work studying instantons with continuous symmetries, we do not have a compact gauge group. However, we are able to apply the same methods as in the proof of Theorem 1.1. This is because on the subset of the space $\mathcal{X}$ corresponding to instantons, we are able to reduce the gauge group to a compact subgroup [Lan24b, §3].
@@
In this section, we construct an example and use Proposition 1.2 to find symmetric points. Previous work has used Theorem 1.1 and Proposition 1.2 to study symmetric gauge theoretic objects in more involved case studies [CDL+22, Lan24a, Lan24b].

### v2

@@ 1 Introduction
The aforementioned examples were found using the relationships between the various topological solitons and related objects. For instance, the examples of instantons and hyperbolic monopoles were found using their corresponding ADHM data—quaternionic matrices satisfying non-linear constraints. The topological solitons themselves do not fit into the framework studied in this paper. Indeed, the space of objects and the space of gauge transformations are infinite-dimensional. However, the related objects do fit in this framework and previous work of mine has used these related objects to identify novel examples of instantons, hyperbolic monopoles, and Euclidean monopoles [CDL+22, Lan24a, Lan24b, Lan25].
@@
This work differs from the above in that we examine general Lie group actions instead of finite, abelian, or absolute Galois group actions. This choice of action allows us to study different kinds of symmetries. Moreover, by choosing Lie group actions, we can make use of smoothness to differentiate the equations of symmetry, providing a linear constraint, which greatly simplifies any potential non-linear constraints. Indeed, my previous work using these results have identified many novel topological solitons [CDL+22, Lan24a, Lan24b, Lan25]. Additionally, this work differs as we do not restrict ourselves to a specific moduli space and group action.
In Section 2, we prove Theorem 1.1, which deals with compact Lie groups and first appeared in my thesis [Lan24b, Theorem 1.1.1 & Appendix A]. As compact Lie groups can be realized as matrix Lie groups, the Lie algebra homomorphisms give us representations of Lie algebras, so we reduce the problem to the realm of representation theory. We also prove Proposition 1.2, which allows us to differentiate the equations of symmetry when the symmetry group is one-dimensional but not necessarily compact. In Section 2.1, we use Proposition 1.2 to find symmetric points in an example. For in depth examples of the use of this result, see my thesis as well as my previous papers, where I identify novel examples of symmetric instantons as well as hyperbolic and Euclidean monopoles [CDL+22, Lan24a, Lan24b, Lan25]. In Section 2.2, we show that the group of symmetries of an object is a closed Lie subgroup, which helps narrow down the kinds of symmetries that we need to study when looking at objects with differing continuous symmetries.
@@ 2 Fixed points
In some cases, we can get even more information about fixed points using representation theory. Suppose $\mathcal{X}$ is a vector space and $\mathcal{G}$ as well as $\mathcal{S}$ act linearly on $\mathcal{X}$, so that our actions are representations. For a symmetric point $A\in\mathcal{X}$, Theorem 1.1 tells us that $\mathrm{Lie}(\mathcal{S})$ acts on $A$ in two ways: through the action of $\mathrm{Lie}(\mathcal{S})$ directly and through the action of $\mathrm{Lie}(\mathcal{G})$ via $\rho$. Hence, there are some representations $(V_{1},\lambda_{1})$ and $(V_{2},\lambda_{2})$ of $\mathrm{Lie}(\mathcal{S})$ such that $A$ belongs to a trivial component of $(V_{1},\lambda_{1})\otimes(V_{2},\lambda_{2})$. This not only helps us determine the dimension of the vector space of solutions (the number of trivial components), but also tells us what representations $\rho$ to consider, as the tensor product must contain trivial components in order to have non-zero symmetric objects and the tensor product depends on $\rho$.
Theorem 1.1 requires compact groups. However, hope is not necessarily lost when you do not have compactness. Indeed, in previous work studying instantons with continuous symmetries, we do not have a compact gauge group. However, we are able to apply the same methods as in the proof of Theorem 1.1. This is because on the subset of the space $\mathcal{X}$ corresponding to instantons, we are able to reduce the gauge group to a compact subgroup [Lan25].
@@
In this section, we construct an example and use Proposition 1.2 to find symmetric points. Previous work has used Theorem 1.1 and Proposition 1.2 to study symmetric gauge theoretic objects in more involved case studies [CDL+22, Lan24a, Lan24b, Lan25].
@@ 2.2 Closed sets
Proposition 2.4 requires a proper action. However, proving that an action is proper is challenging. It is much simpler when the group is compact. However, just as before, hope is not necessarily lost when you do not have compactness. Indeed, in previous work studying instantons with continuous symmetries, we do not have a compact gauge group. However, we are able to apply the same methods as above, when $\mathcal{G}$ is compact [Lan25, Proposition 3.4].